# Will Neural Scaling Laws Activate Jevons’ Paradox in AI Labor Markets? A Time-Varying Elasticity of Substitution (VES) Analysis

### v1

@@ Abstract
AI industry leaders often use the term “Jevons’ Paradox.” We explore the significance of this term for artificial intelligence adoption through a time-varying elasticity of substitution framework. We develop a model connecting AI development to labor substitution through four key mechanisms: (1) increased effective computational capacity from both hardware and algorithmic improvements; (2) AI capabilities that rise logarithmically with computation following established neural scaling laws; (3) declining marginal computational costs leading to lower AI prices through competitive pressure; and (4) a resulting increase in the elasticity of substitution between AI and human labor over time. Our time-varying elasticity of substitution (VES) framework, incorporating the Gørtz identity, yields analytical conditions for market transformation dynamics. This work provides a simple framework to help assess the economic reasoning behind industry claims that AI will increasingly substitute for human labor across diverse economic sectors.
@@ 1 Introduction
Many industry leaders have appealed to Jevons’ Paradox (1866) where price declines lead to more usage when talking about the progress of AI. Also, industry leaders have stated that the AI scaling laws will lead to quality improvements (Kaplan et al., 2020; Hoffmann et al., 2022) that will also lead to more usage. [^1] For example, Sam Altman (2025), in his blog entitled “Three Observations,” stated that “1. The intelligence of an AI model roughly equals the log of the resources used to train and run it. 2. The cost to use a given level of AI falls about 10X every 12 months, and lower prices lead to much more use. 3. The socioeconomic value of linearly increasing intelligence is super-exponential in nature.” The first point restates the AI scaling law while the second point is consistent with Jevons’ Paradox. Satya Nadella (2025) stated that “Jevons paradox strikes again! As AI gets more efficient and accessible, we will see its use skyrocket, turning it into a commodity we just can’t get enough of.” Driving much of this has been the decreased marginal cost of computation and competition that links the lower marginal costs to the lower cost of AI. [^2] Ho et al. (2024) report that the “… compute required to reach a set performance threshold has halved approximately every 8 months, … .” Is Jevons’ Paradox mere industry jargon, or does it provide economically meaningful insights relevant to AI progress?
Jevons’ Paradox occurs when increased efficiency in a resource leads to greater total demand due to both intensive margins (existing users consuming more) and extensive margins (new users or applications entering the market which may take time). For example, the extensive margin of uses from electricity has dwarfed the initial uses of electricity for lighting, but many uses took many years to emerge. In elasticity terms, this implies a long-term price elasticity of demand less than $-1$ (i.e., highly elastic, $E<-1$), where the percentage increase in quantity demanded exceeds the percentage decrease in price. The AI context involves both pecuniary price decreases as well as quality increases for a reduction in the quality-adjusted price.
@@
This analysis develops a framework that incorporates these statements so as to make the underlying logic clearer. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework, drawing on the identity presented in Gørtz (1977). [^3] Smith (1979) points out that the identity was not necessarily new, but implied by Hicks and Allen (1934), although the derivation was new. We demonstrate conditions under which complete market transformation occurs, highlighting the role of both price elasticities and substitution elasticities.
@@
To address the question of Jevons’ Paradox in the context of AI competing with human labor, and whether the quantity demanded of an AI-provided good rises relatively more than its price decreases, we analyze this using a CES utility function (Arrow et al., 1961) and an identity between price elasticity and the elasticity of substitution from Gørtz (1977).
@@
We further explore the effects of exponential declines in the marginal cost of effective computation which via competition leads to exponential declines in the price of AI-provided goods and services as well as the increased quality of AI goods and services which flows from the increased effective computation created by the lower marginal costs via the AI “scaling laws.” The combination of exponential declines in the price of AI-provided goods and services and increases in quality (increased ability to substitute) leads to AI adoption with an early phase dominated by price making way for a mature phase that uses quality increases for market domination. In fact, unless AI quality rises sufficiently (elasticity of substitution above 1), no level of AI price declines (provided $p_{A}>0$) will result in complete capture of markets.
This paper makes three primary contributions: First, we formalize the connection between elasticity of substitution and own-price elasticities in AI-human substitution using the Gørtz (1977) identity. Second, we derive analytical conditions for market transformation using a variable elasticity of substitution specification. Third, as an aid to categorization, we set out a five-phase adoption process with distinct economic mechanisms governing each phase that highlights when Jevons’ Paradox occurs.
@@ 2 Jevons’ Paradox and Elasticity
Assume a CES utility function for consumers choosing between human-produced (H) and AI-produced (A) goods in (1),
$U$ $=(H^{\rho}+A^{\rho})^{1/\rho}$ (1)
where $\rho=\frac{\sigma-1}{\sigma}$ and $\sigma$ represents the elasticity of substitution. Gørtz (1977) presented an identity in (2) that relates the own price elasticities and the elasticity of substitution,
$\sigma+E_{AA}+E_{HH}=-1$ (2)
where for this problem $E_{AA}$ is the own-price elasticity of AI goods, and $E_{HH}$ is the own-price elasticity of human goods. The Gørtz identity provides an analytical link between the elasticity of substitution and own-price elasticities in this AI-human substitution framework. This relation allows characterization of market transitions more completely than examining either elasticity measure in isolation. For AI-produced goods, the elasticity identity reveals boundary conditions for market dominance that depend on both the intrinsic substitutability between AI and human production and the market’s response to price changes for each good, service, or task.
@@
Given normalized human prices ($p_{H}=1$) and AI prices $p_{A}\in(0,1)$, the AI expenditure or revenue share, $r_{A}$, appears in (3). [^4] The ratio of the AI price to human prices is not that easy to quantify. First, AI prices involve the cost of the tokens for input and output. Second, queries differ across numbers of tokens required. Third, many queries also involve a human and so the AI cost is really a hybrid AI-human cost. Fourth, many tasks may require multiple queries. Finally, the cost of a human query also varies by task. Therefore, the ratio of the AI price to the human price is not obvious in magnitude, but should be less than 1.
$r_{A}$ $=\frac{1}{1+p_{A}^{\,\sigma-1}}$ (3)
What happens as the price of AI declines? Assume $p_{A}\to 0$,
$r_{A}$ $\to 1,\text{ for }\sigma>1$ (4)
@@
This reveals that no matter what the price of AI, if AI does not have sufficient quality to substitute for human products, services, or tasks, it cannot achieve market domination in expenditures $r_{A}$.
@@
It may aid understanding of the role of price and the elasticity of substitution to consider the elasticity of the revenue share $r_{A}$ to the AI price $p_{A}$ in (5).
$E_{r_{A},p_{A}}=-\frac{(\sigma-1)p_{A}^{\sigma-1}}{1+p_{A}^{\sigma-1}}$ (5)
Clearly, the revenue rises as price decreases when $\sigma>1$. However, for $\sigma\in(1,2)$, a concave relation exists for $p_{A}^{\sigma-1}$ and this becomes convex for $\sigma>2$. This arises because for $p_{A}<1$, $\sigma\in(1,2)$ moves $p_{A}^{\sigma-1}$ towards 1, while $\sigma>2$, moves $p_{A}^{\sigma-1}$ towards 0.
For the revenue share $r_{A}$ this aligns with Gørtz’s identity since one can substitute out $E_{HH}$ to arrive at (6).
$E_{AA}$ $=-\sigma+(\sigma-1)r_{A}$ (6)
Equation (6) shows that as the AI market share approaches complete dominance ($r_{A}\to 1$), the own-price elasticity converges to unity ($E_{AA}\to-1$), representing the cessation of Jevons’ Paradox. This transition from elastic to unitary elastic demand signifies market saturation, where further price reductions no longer generate proportionally larger quantity increases. The convergence occurs precisely when substitution opportunities are exhausted, representing a natural boundary condition in the AI adoption process.
@@ 3 Exponential Declines in the AI Price, Scaling Laws, and Increased Elasticity of Substitution (VES)
Specifically, assume that improvements in AI capabilities increase the elasticity of substitution by means of the scaling laws which show a linear improvement in AI performance as inputs (in this case computation) increase logarithmically as in (8). Note, the actual value of $\delta>0$ could differ across goods, services, tasks, contexts, or organizations.
@@
The compound effect parameter $\phi$ captures two forces driving AI adoption. Namely, the rate at which computation increases ($g$) and the sensitivity of the elasticity of substitution $\delta$ to computational improvements. In other words, $\phi$ specifies the elasticity of substitution growth rate. This form links the technical AI scaling laws literature with economic substitution theory. Note, the models improve in terms of loss with repect to inputs at different reported rates, Kaplan et al. (2020) reported a rate of loss decline of around $0.05$. So if this directly translated into $\delta$, one could entertain a prior on $\delta$ of around $0.05$. However, as mentioned above, this likely varies substantially across various uses. [^6] We use the word “prior” in conjunction with the parameters $g$, $\delta$ to convey a sense of uncertainty as to their values. The model treats the parameters as constants. One could certainly place an informative prior on these parameters and simulate across the potential outcomes which might yield a more dispersed, non-symmetric distribution. We do not pursue this further as the goal is create a simple framework while qualitatively acknowledging uncertainty.
Consistent with Ho et al. (2024), assume exponentially declining AI prices in (10),
@@
where $10<p_{A}(0)<$ 1 represents the initial price ratio between AI and human goods. [^7] Although we model the declining marginal cost cost of computation and the increased quantity of computation separately, they obviously are related through the demand equation. However, there are other shifters of demand and so we will retain the simplicity of separate equations. In addition, to the degree that AI work, especially when it has a low elasticity of substitution for human work, in reality is hybrid AI-human work. Therefore, pecuniary AI price declines may overstate the price declines in actual tasks.
@@
Substituting this modified price function into equation (3) while maintaining the variable elasticity of substitution from equation (8), leads to a generalized logistic function (11) that represents the revenue share $r_{A}(t)$ based on parameter values for $a,b,c$ in (12).
$r_{A}(t)$ $=\frac{1}{1+\exp\Bigl{(}a+b\cdot t+c\cdot t^{2}\Bigr{)}}$ (11)
$a$ $=-\ln\bigl{(}p_{A}(0)\bigr{)},\quad b=d+\phi\cdot\ln\bigl{(}p_{A}(0)\bigr{)},\quad c=-d\cdot\phi$ (12)
The revenue share $r_{A}(t)$ will monotonically increase over time after time $t_{0}$ (where $t_{0}$ could be 0, 1, or some other value) provided the initial price is low enough as shown in (13).
$p_{A}(0)<\exp\left(d\cdot\left(2\cdot t_{0}-\frac{1}{\phi}\right)\right)$ (13)
This comes from taking the derivative of the argument in the exponential in (11) ($b+2\cdot c\cdot t$) and making sure that it is strictly negative for $t\geq t_{0}$.

### v2

@@ Abstract
We develop a formal economic framework to analyze whether neural scaling laws in artificial intelligence will activate Jevons’ Paradox in labor markets, potentially leading to increased AI adoption and human labor substitution. By using a time-varying elasticity of substitution (VES) approach, we establish analytical conditions under which AI systems transition from complementing to substituting for human labor. Our model formalizes four interconnected mechanisms: (1) exponential growth in computational capacity ($C(t)=C(0)\cdot e^{g\cdot t}$); (2) logarithmic scaling of AI capabilities with computation ($\sigma(t)=\delta\cdot\ln(C(t)/C(0))$); (3) declining AI prices ($p_{A}(t)=p_{A}(0)\cdot e^{-d\cdot t}$); and (4) a resulting compound effect parameter ($\phi=\delta\cdot g$) that governs market transformation dynamics. We identify five distinct phases of AI market penetration, demonstrating that complete market transformation requires the elasticity of substitution to exceed unity ($\sigma>1$), with the timing determined primarily by the compound parameter $\phi$ rather than price competition alone. These findings provide an analytical framing for evaluating industry claims about AI substitution effects, especially on the role of quality versus price in the technological transition.
@@ 1 Introduction
Many industry leaders appeal to Jevons’ Paradox (1866) – the principle that efficiency gains that lower the price of a resource can increase its overall consumption – when discussing AI progress. Furthermore, industry leaders state that AI scaling laws that deliver quality improvements with increased computation (Kaplan et al., 2020; Hoffmann et al., 2022) will drive usage. [^1] For example, Sam Altman (2025), in his blog entitled “Three Observations,” stated that “1. The intelligence of an AI model roughly equals the log of the resources used to train and run it. 2. The cost to use a given level of AI falls about 10X every 12 months, and lower prices lead to much more use. 3. The socioeconomic value of linearly increasing intelligence is super-exponential in nature.” The first point restates the AI scaling law while the second point is consistent with Jevons’ Paradox. Satya Nadella (2025) stated that “Jevons paradox strikes again! As AI gets more efficient and accessible, we will see its use skyrocket, turning it into a commodity we just can’t get enough of.” Decreasing marginal costs of effective computation (which includes software as well as hardware advances), coupled with competition linking these lower costs to lower AI prices, are seen as key drivers. [^2] Ho et al. (2024) report that the “… compute required to reach a set performance threshold has halved approximately every 8 months … .” Is Jevons’ Paradox mere industry jargon, or does it provide economically meaningful insights relevant to AI progress?
@@
We develop a partial equilibrium framework that incorporates these statements to clarify the underlying logic of AI adoption. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework. We demonstrate conditions under which complete market transformation occurs, highlighting the role of both price elasticities and substitution elasticities.
Jevons’ Paradox occurs when increased efficiency in a resource leads to greater total demand due to both intensive margins (existing users consuming more) and extensive margins (new users or applications entering the market, a process that may take time). For example, the extensive margin of uses from electricity has dwarfed the initial uses of electricity for lighting, but many uses took many years to emerge. In elasticity terms, this implies a long-term price elasticity of demand less than $-1$ (i.e., highly elastic, $E<-1$), where the percentage increase in quantity demanded exceeds the percentage decrease in price.
However, industry statements seem to point to obtaining a larger expenditure share of the entire market from the consumers which translates to revenue share of the entire market for the producers (for that product or service). The tremendous investments made by firms in AI (Tobin and Karaian, 2025; Waters and Bradshaw, 2025) seem to support this grander mission. In this sense, the use of Jevons’ Paradox, a statement about price elasticity, could be viewed as a sleight-of-hand. A more indirect way to signal an AI capable of substantial displacement of human labor, a potentially less popular statement. The goal of this paper is to explore the meaning of Jevons’ Paradox as industry sources use it given the context of the massive investments that have been made in AI. We do not try to predict their success or failure in their endeavors, but to understand the motivation of this ambition.
@@
For simplicity, we begin with a Constant Elasticity of Substitution (CES) utility function (Arrow et al., 1961) to address the question of Jevons’ Paradox in the context of AI competing with human labor, and whether the quantity demanded of an AI-provided good rises relatively more than its price decreases. However, the CES approach holds the elasticity of substitution constant, but a salient aspect of AI is its rapidly evolving capabilities which suggests an elasticity of substitution that rises over time.
@@
Accordingly, we extend our analysis using a Variable Elasticity of Substitution (VES) framework to explore the effects of exponential declines in the marginal cost of effective computation which, via competition and scaling laws, leads to declines in the price and increases in the quality of AI-provided goods and services. The combination of exponential declines in the price of AI-provided goods and services and increases in quality (increased ability to substitute) leads to time-varying AI adoption with an early phase dominated by price, transitioning to a mature phase based on quality increases to achieve market domination. In fact, unless AI quality rises sufficiently (elasticity of substitution above 1), no level of AI price declines (provided $p_{A}>0$) will result in complete capture of markets.
We explore generalizations which allow for slower power law quality improvements and price declines. In addition, we allow for uncertainty in the quality improvements and heterogeneity in the initial elasticity of substitution. Finally, we show how examining differences in log-odds of market shares can simplify possible analysis.
This paper makes four contributions: (1) it translates the concept of Jevons’ Paradox (own-price elasticity of less than $-1$) into the implied elasticity of substitution associated with the idea of rising revenue share in response to lower prices of AI; (2) derives analytical conditions for market transformation under a time-varying elasticity of substitution (VES) framework motivated by the neural scaling laws; (3) generalizes the VES framework to allow for power law quality improvements and price declines as well as allowing for uncertainty in quality improvements and heterogeneity in elasticities of substitution; and (4) proposes a five-phase adoption taxonomy highlighting when Jevons’ Paradox emerges.
Section 2 establishes a static CES framework to explore Jevons’ Paradox, while Section 3 introduces a dynamic VES model to capture AI’s evolving role with generalizations. Section 4 uses specific parameters to depict some possible trajectories. The conclusion in Section 5 ties these insights together and outlines future research directions.
@@ 2 Jevons’ Paradox and Elasticity
The constant elasticity of substitution (CES) utility function provides a means of studying how the elasticity of substitution affects the trade-off between consumer choices. We begin with the weighted CES utility function for consumers choosing between human-produced (H) and AI-produced (A) goods in (1),
$U$ $=\bigl{(}(1-\alpha)\cdot H^{\rho}+\alpha\cdot A^{\rho}\bigr{)}^{1/\rho}$ (1)
where $\alpha\in(0,1)$ represents the relative preference weight for AI-produced goods over human-produced goods, $\rho=\frac{\sigma-1}{\sigma}$, and $\sigma$ represents the elasticity of substitution. Note, we later allow the elasticity of substitution to vary over time ($\sigma(t)$) which would lead to a variable elasticity of substitution utility function or VES. However, we first wish to develop the simpler CES case.
@@
Given normalized human prices ($p_{H}=1$) and AI prices $p_{A}\in(0,1)$, the AI expenditure or revenue share, $r_{A}$, appears in (2). [^3] The ratio of the AI price to human prices is not that easy to quantify. First, AI prices involve the cost of the tokens for input and output. Second, queries differ across numbers of tokens required. Third, many queries also involve a human and so the AI cost is really a hybrid AI-human cost. Fourth, many tasks may require multiple queries. Finally, the cost of a human query also varies by task. Therefore, the ratio of the AI price to the human price is not obvious in magnitude.
$r_{A}$ $=\frac{\alpha\cdot p_{A}^{1-\sigma}}{(1-\alpha)+\alpha\cdot p_{A}^{1-\sigma}}=\frac{1}{1+\displaystyle\frac{1-\alpha}{\alpha}\cdot p_{A}^{\sigma-1}}$ (2)
We focus on expenditure share from the consumers which translates to revenue share for the producers (for that product or service) as the industry statements as well as tremendous investments seem to aim for market domination.
@@
What happens to the revenue share as the price of AI declines to low levels? If the elasticity of substitution $\sigma$ exceeds 1, the revenue share $r_{A}$ approaches 1. If $\sigma=1$, we have $r_{A}=\alpha$ for any AI price. If $\sigma<1$, then for every strictly positive price with $10<p_{A}<$ 1 we obtain $r_{A}<\alpha$; a fortiori, as $p_{A}\to 0$ we have $r_{A}\to 0$. This reveals that no matter how low the AI price falls, if AI does not have sufficient quality to substitute for human products, services, or tasks, it cannot achieve market domination in terms of revenue share $r_{A}$.
@@
To aid understanding of the role of price and the elasticity of substitution consider the elasticity of the revenue share $r_{A}$ to the AI price $p_{A}$ in (3).
$E_{r_{A},p_{A}}$ $=\frac{\partial\ln r_{A}}{\partial\ln p_{A}}=-(\sigma-1)\cdot(1-r_{A})$ (3)
Clearly, the revenue rises as price decreases when $\sigma>1$ and falls when $\sigma<1$.
Although the CES model quantifies the revenue market share as a function of price for a given level of elasticity of substitution $\sigma$, it seems restrictive in that if $\sigma\leq 1$, $r_{A}\leq\alpha$ and low levels of $p_{A}$ cannot break through this level. Insofar as the models have continued to rise in performance, as measured by benchmarks, this suggests that a model needs to allow $\sigma$ to evolve to reflect the greater model capabilities. To address this, we adopt a variable elasticity of substitution (VES) model in Section 3.
@@ 3 Exponential Declines in the AI Price, Scaling Laws, and Increased Elasticity of Substitution (VES)
Specifically, assume that improvements in AI capabilities increase the elasticity of substitution by means of the scaling laws which show a linear improvement in AI performance (lower loss) as inputs (in this case computation) increase logarithmically as in (5). Although one could propose other functional forms besides the proportional specification in (5), we select this specification for simplicity and because it may approximate a more elaborate specification. Note, the actual value of $\delta>0$ could differ across goods, services, tasks, contexts, or organizations.
@@
The compound effect parameter $\phi$ captures two forces driving AI adoption. Namely, the rate at which computation increases ($g$) and the sensitivity of the elasticity of substitution $\delta$ to computational improvements. In other words, $\phi$ specifies the elasticity of substitution growth rate. This form links the technical AI scaling laws literature with economic substitution theory. Note, the models improve in terms of loss with respect to inputs at different reported rates. Kaplan et al. (2020) reported a rate of loss decline of around $0.05$. So if this directly translated into $\delta$, one could entertain a prior on $\delta$ of around $0.05$. However, as mentioned above, this likely varies substantially across various uses. [^5] We use the word “prior” in conjunction with the parameters $g$, $\delta$ to convey a sense of uncertainty as to their values. The model treats the parameters as constants. One could certainly place an informative prior on these parameters and simulate across the potential outcomes which might yield a more dispersed, non-symmetric distribution. We do not pursue this further as the goal is create a simple framework while qualitatively acknowledging uncertainty. Note, at small, strictly positive $t$ the AI and human shares are almost fixed as $\sigma$ is near 0 and this implies almost perfect complements. However, as $t$ rises $\sigma$ becomes increasingly positive and this artifact vanishes.
The simple form of (6) means that one can solve for the time when $\sigma=1$, denoted at $t^{*}$ in (7).
$t^{*}$ $=\dfrac{1}{\phi}$ (7)
Moreover, the time for any particular value of $\sigma$ is linear and so $\sigma=2$ implies $t=2\cdot t^{*}$ and so forth.
The previous equations dealt with improvements in AI quality leading to greater elasticities of substitution. However, AI pricing affects adoption and producer revenue. Consistent with Ho et al. (2024), assume exponentially declining AI prices in (8),
@@
where $10<p_{A}(0)<$ 1 represents the initial price ratio between AI and human goods. [^6] Although we model the declining marginal cost of computation and the increased quantity of computation separately, they obviously are related through the demand equation. However, there are other shifters of demand and so we will retain the simplicity of separate equations. In addition, to the degree that AI work, especially when it has a low elasticity of substitution for human work, in reality is hybrid AI-human work. Therefore, pecuniary AI price declines may overstate the price declines in actual tasks.
@@
Substituting this modified price function (8) into equation (2) while maintaining the variable elasticity of substitution from equation (5), leads to a generalized logistic function (9) that represents the revenue share $r_{A}(t)$ based on parameter values for $a,b,c$.
$r_{A}(t)$ $=\dfrac{1}{1+\exp\bigl{(}a+b\cdot t+c\cdot t^{2}\bigr{)}}$ (9)
$a$ $=\ln\biggl{(}\frac{1-\alpha}{\alpha}\biggr{)}-\ln p_{A}(0),\ b=d+\phi\cdot\ln p_{A}(0),\ c=-d\cdot\phi$
Arguably, $c$ represents the most important term in (9). First, it captures three of the most important influences on AI competitiveness: $d$ the rate of price declines; $g$ the growth rate in computational capacity; and $\delta$ the growth rate in quality with respect to the rate of computational growth. Since $d,g,\delta>0$ and $c=-(d\cdot g\cdot\delta)$, then $c<0$. Second, because $c<0$, $t^{2}$ will dominate the other terms for large $t$ and thus the magnitude of $c$ determines the speed of the later part of the transition. Finally, the additional quadratic term of $c\cdot t^{2}$ allows acceleration beyond the standard logistic’s symmetrical ’S’ shape. This acceleration could make the AI transition occur abruptly.
Although the revenue share function for a given level of $\phi$, $r_{A}(t|\phi)$, provides direct insight into market penetration, transforming it can reveal underlying linearities and simplify comparisons between different scenarios. Specifically, the logit transformation, defined as $\operatorname{logit}(p)=\ln(p/(1-p))$, proves particularly useful.
3.1 Dynamics under Power-Law Quality Improvement and Price Declines with Random $\sigma_{0}$ and $\phi$
In this sub-section we explore generalizations to the deterministic model developed above. Specifically we (1) allow for possible heterogeneity in the elasticity of substitution or the uncertainty of the quality improvement rate parameter $\phi$; (2) for power law growth in $\sigma(t)$ and in price declines $p_{A}(t)$; and (3) show how differencing the log-odds of revenue share for different tasks reduces the complexity of the relations.
We begin by adopting the power law forms for the elasticity of substitution over time in (10) and in price declines in (11) using the definition $\beta_{0}$ in (12).
$\sigma(t)$ $=\sigma_{0}+\phi\cdot t^{k}$ (10)
$p_{A}(t)$ $=p_{A}(0)\cdot\exp\left(-d\cdot t^{\xi}\right)$ (11)
$\beta_{0}$ $=\ln p_{A}(0)$ (12)
This captures the idea that quality improvements and price decline rates may diminish over time, assuming $k,\xi\in(0,1)$. Now the revenue share $r_{A}(t)$ is given by (13) in terms of $X(t)$ defined in (14), the complement of which equals the logit of the revenue share over time in (15).
$r_{A}(t)$ $=\left(1+\exp\left(X(t)\right)\right)^{-1}$ (13)
$X(t)$ $=\ln\left(\left(1-r_{A}(t)\right)/r_{A}(t)\right)$ (14)
$\operatorname{logit}r_{A}(t)$ $=-X(t)$ (15)
One can rewrite $X(t)$ in (16) as a linear combination of the baseline elasticity parameter $\sigma_{0}$ and the quality improvement parameter $\phi$ using the definitions for $\Theta(t)$ in (18), $\Upsilon(t)$ in (17), and $\Xi(t)$ in (19)
$X(t)$ $=\Theta(t)+\Upsilon(t)\cdot\sigma_{0}+\Xi(t)\cdot\phi$ (16)
$\Upsilon(t)$ $=\beta_{0}-d\cdot t^{\xi}$ (17)
$\Theta(t)$ $=\ln\left(\frac{1-\alpha}{\alpha}\right)-\Upsilon(t)$ (18)
$\Xi(t)$ $=t^{k}\cdot\Upsilon(t)$ (19)
This provides the apparatus for examining the generalizations to follow.
3.1.1 Task elasticity of substitution and quality improvement heterogeneity and uncertainty
Because equation (16) is linear in $\sigma_{0}$ and $\phi$, this facilitates letting these be random variables instead of constant parameters as in (20).
$\tilde{X}(t)$ $=\Theta(t)+\Upsilon(t)\cdot\tilde{\sigma}_{0}+\Xi(t)\cdot\tilde{\phi}$ (20)
This allows modeling some task heterogeneity by perhaps assuming a exponential distribution where some tasks are inherently easy to substitute while a great number start off near 0. It also allows for modeling uncertainty in the distribution for $\tilde{\phi}$. However, regardless of the distribution, the expectation of $X(t)$ in (21) is linear in the individual expectations of $\tilde{\sigma}_{0}$ and $\tilde{\phi}$.
$\mathbb{E}\left[X(t)\right]$ $=\Theta(t)+\Upsilon(t)\cdot\mathbb{E}\left[\sigma_{0}\right]+\Xi(t)\cdot\mathbb{E}\left[\phi\right].$ (21)
If we assume $\phi$ is common across tasks (i.e., non-random or its mean is used) and only $\sigma_{0}$ is random, (22) shows the variance of $X(t)$.
$\operatorname{Var}\left[X(t)\right]$ $=\Upsilon(t)^{2}\cdot\operatorname{Var}[\sigma_{0}].$ (22)
3.1.2 Power law quality improvements and price declines
In this more general framework, the long-run behavior of the AI revenue share $r_{A}(t)$ depends on the limiting behavior of $X(t)$. We further expand $X(t)$ in (24) and repeat the definition of $\Upsilon$ in (23). We solve for $\sigma(t)=1$ as shown in (25) formed by substituting $\sigma_{t}=1$ into (10) and solving for $t_{G}^{*}$, the generalized timeline. At that point, $(t^{*}_{G})^{k}\cdot\phi=(1-\sigma_{0})$ and so the term in (24) associated with $\Upsilon(t)$, $(\sigma_{0}-1+t^{k}\cdot\phi)=0$. For $t>t_{G}^{*}$, $(\sigma_{0}-1+t^{k}\cdot\phi)>0$. Since $\beta_{0}<0$, $\Upsilon(t)<0$ and becomes more negative with increasing $t$. Although for $\alpha<0.5$, $\ln\left(\frac{1-\alpha}{\alpha}\right)>0$, this stays constant over time. Therefore, for sufficiently large $t$, $X(t)<0$ and becomes progressively more negative with $t$, leading to a greater AI market share.
$\Upsilon(t)$ $=\beta_{0}-d\cdot t^{\xi}$ (23)
$X(t)$ $=\ln\left(\frac{1-\alpha}{\alpha}\right)+\Upsilon(t)\cdot(\sigma_{0}-1+t^{k}\cdot\phi)$ (24)
$t_{G}^{*}$ $=\left(\dfrac{1-\sigma_{0}}{\phi}\right)^{1/k}$ (25)
However, the rise in market share could become slow, depending on the values of $k,\phi$ given the nature of power laws. For example, if $k=0.5$, it would cause the timeline to square, all else equal, and so 10 years if $k=1$ becomes 100 years when $k=0.5$. If quality improvements grow slowly $\phi$ becomes small which also increases the timeline, perhaps dramatically. Note, allowing for $\tilde{\sigma}_{0}$ and $\tilde{\phi}$ in (25) is possible, which would create $\tilde{t}_{G}^{*}$ which could have a complicated distribution given that it is the ratio of random variables.
3.1.3 Simplification from differencing revenue share log-odds across tasks
As a final generalization, consider two tasks, indexed $j\in\{1,2\}$, differing only in their baseline elasticities $\sigma_{0,1}$ and $\sigma_{0,2}$ while sharing the same parameters $\alpha,\beta_{0},d,k,\xi,$ and $\phi$. Let $X_{j}(t)$ be the exponent term for task $j$. The difference is appears in (26).
$\Delta_{X}(t)$ $=X_{1}(t)-X_{2}(t)$
$=\left(\Theta(t)+\Upsilon(t)\cdot\sigma_{0,1}+\Xi(t)\cdot\phi\right)-\left(\Theta(t)+\Upsilon(t)\cdot\sigma_{0,2}+\Xi(t)\cdot\phi\right)$
$=\Upsilon(t)\cdot\left(\sigma_{0,1}-\sigma_{0,2}\right).$ (26)
The relative AI revenue shares of the two tasks are driven by $\Delta_{X}(t)$. This differences out $\Theta(t)$ and $\Xi(t)\cdot\phi$ which simplifies the analysis.
3.2 Five Phase Categorization